# UNGER: Generative Recommendation with A Unified Code via Semantic and Collaborative Integration

### v2

@@ 3.3. Stage I: Item Unicodes Generation
Unified Codes. After the training is finished, following previous work (Wang et al., 2024b), we apply the widely-used hierarchical k-means algorithm to the integrated embeddings for quantization. In this process, the embeddings are initially partitioned into $K$ clusters. Each cluster is then recursively subdivided into K child clusters until each child cluster contains a single item. Finally, for each item, the cluster indices at each hierarchical level are recorded sequentially, forming a list that serves as the item’s unicode.
@@ 3.4. Stage II: Generative Recommendation
Encoding process. On the encoder side, given a user’s interaction history $X=[x_{1},x_{2},\dots,x_{t-1}]$, we input it into an encoder consisting of stacked multi-head self-attention layers and feed-forward layers, following the Transformer architecture. The encoder processes $X$ to produce a feature representation $H$, which captures the user’s interests and will be passed to the decoder.
Decoding process. On the decoder side, items are mapped to their unicodes derived in Stage I. Our objective changes from predicting the next item $x_{t}$ to predicting its item unicode $[c_{1},c_{2},\dots,c_{L}]$. For training, a special token <BOS> is prepended to the item unicode to construct the decoder input. The generative recommendation loss is computed using a cross-entropy loss function, defined as follows:
@@
Intra-modality knowledge distillation task. To compensate for the information loss caused by the quantization process in the first stage, we introduce an intra-modality knowledge distillation task to distill the global knowledge. Inspired by the use of the [CLS] token in BERT for capturing global context in classification tasks, we append a learnable token $[c_{\text{dis}}]$ to the end of the decoder input, allowing gradients to propagate through each preceding unicode. This design encourages $[c_{\text{dis}}]$ to capture global information.
Specifically, the final layer output corresponding to $[c_{\text{dis}}]$ is used in a global contrastive learning objective over the item corpus: the positive sample is the integrated embedding $E_{t}$ of item $x_{t}$ learned from the first stage, and the negative sample is a randomly selected integrated embedding $E_{\text{neg}}$ from other items in the corpus, excluding $x_{t}$. This objective pulls the $[c_{\text{dis}}]$ final layer output closer to the positive sample $E_{t}$ and pushes it away from the negative sample $E_{\text{neg}}$. The loss function for the distillation task is defined as follows:
@@
where $c_{\text{dis}}$ denotes the final layer output of the decoder for the special token $[c_{\text{dis}}]$; $E_{\text{neg}}$ is the integrated embedding for a negative sample as learned in Stage I, while $E_{t}$ represents the integrated embedding of the target item, also learned in Stage I.
Joint optimization. Finally, in Stage II, we jointly optimize these two loss functions. The overall loss function for Stage II is defined as follows:
@@ 3.5. Training and Inference
Training. we employ a two-stage training process. In the first stage, we use DIN as the backbone and optimize the model with loss $\mathcal{L}_{\text{stage I}}$. After the first stage training is complete, we extract the item integrated embeddings and derive item unicodes. In the second stage, we use a Transformer for generative recommendation, optimizing it with loss $\mathcal{L}_{\text{stage II}}$.
Inference. During inference phase, we rely exclusively on the Transformer model. In the decoding process, beam search is applied iteratively to generate each token within the item unicode sequence. Once finished, the item unicodes are mapped to their corresponding items through the item-unicode lookup table, creating a ranked recommendation list based on confidence scores to produce the final top-$k$ results.
@@ 3.6. Computational and Storage Costs Analysis
In our approach, we employ a unified code to encode both semantic and collaborative information. Regarding computational costs, the encoder processes the user interaction sequence only once, leaving the main computational burden to the decoder, which iteratively generates each token of the item code. Assuming the code length is fixed and decoding an item code requires $\mathcal{O}(n)$ time, existing methods (Wang et al., 2024b; Kim et al., 2024) that use separate codes for each modality require $K$ parallel decoders to predict codes for all modalities. Furthermore, these methods involve a confidence-based ranking step to derive the final recommendation. Excluding the computational cost of this ranking, their decoding cost is $\mathcal{O}(Kn)$, where $K$ is the number of modalities. In contrast, our method, with its single unified code, achieves a decoding cost of only $\mathcal{O}(n)$.
For storage costs, assuming the storage requirement for one code is $\mathcal{O}(m)$, existing methods require $\mathcal{O}(Km)$ space to store codes for all modalities. By contrast, our method requires only $\mathcal{O}(m)$ space to store multi-modal information within a single unified code.
@@
Table 3. Performance comparison of different methods. The best performance is highlighted in bold while the second best performance is underlined. All the results of UNGER are statistically significant with $p<0.05$ compared to the best baseline models.
Dataset Metric Classical Generative UNGER (Ours)
Beauty Recall@10 0.0283 0.0347 0.0605 0.0347 0.0512 0.0407 0.0647 0.0664 $0.0617^{*}$ $0.0524^{*}$ 0.0836 0.0939
Recall@20 0.0479 0.0556 0.0902 0.0599 0.0773 0.0656 0.0994 0.0915 $0.0924^{*}$ $0.0807^{*}$ 0.1124 0.1289
NDCG@10 0.0137 0.0176 0.0318 0.0170 0.0266 0.0208 0.0327 0.0400 $0.0339^{*}$ $0.0263^{*}$ 0.0525 0.0559
NDCG@20 0.0187 0.0229 0.0394 0.0233 0.0332 0.0270 0.0414 0.0464 $0.0417^{*}$ $0.0335^{*}$ 0.0599 0.0646
Sports Recall@10 0.0204 0.0194 0.0350 0.0191 0.0313 0.0288 0.0385 0.0247 $0.0376^{*}$ $0.0348^{*}$ 0.0441 0.0471
Recall@20 0.0333 0.0314 0.0507 0.0315 0.0477 0.0463 0.0607 0.0375 $0.0577^{*}$ $0.0533^{*}$ 0.0659 0.0710
NDCG@10 0.0110 0.0097 0.0192 0.0099 0.0159 0.0156 0.0204 0.0133 $0.0196^{*}$ $0.0179^{*}$ 0.0236 0.0259
NDCG@20 0.0142 0.0126 0.0231 0.0130 0.0201 0.0200 0.0260 0.0164 $0.0246^{*}$ $0.0226^{*}$ 0.0291 0.0319
Toys Recall@10 0.0176 0.0270 0.0675 0.0203 0.0497 0.0381 0.0700 0.0383 $0.0578^{*}$ $0.0474^{*}$ 0.0714 0.0822
Recall@20 0.0301 0.0420 0.0941 0.0358 0.0716 0.0632 0.1065 0.0483 $0.0838^{*}$ $0.0704^{*}$ 0.1024 0.1154
NDCG@10 0.0084 0.0141 0.0374 0.0099 0.0277 0.0189 0.0376 0.0285 $0.0321^{*}$ $0.0242^{*}$ 0.0505 0.0489
NDCG@20 0.0116 0.0179 0.0441 0.0138 0.0332 0.0252 0.0468 0.0310 $0.0386^{*}$ $0.0300^{*}$ 0.0538 0.0573
@@ 4.1. Experimental Setting
Dataset. We conduct experiments on three public benchmarks commonly used in the sequential recommendation task. For all datasets, we group the interaction records by users and sort them by the interaction timestamps ascendingly. Following (Rendle et al., 2010; Zhang et al., 2019), we only keep the 5-core dataset, which filters unpopular items and inactive users with fewer than five interaction records. Statistics of these datasets are shown in Table 2.
@@
Evaluation Metrics. We employ two broadly used criteria for the matching phase, i.e., Recall and Normalized Discounted Cumulative Gain (NDCG). We report metrics computed on the top 10/20 recommended candidates. Following the standard evaluation protocol (Kang and McAuley, 2018), we use the leave-one-out strategy for evaluation. For each item sequence, the last item is used for testing, the item before the last is used for validation, and the rest is used for training. During training, we limit the number of items in a user’s history to 20.
@@
Implementation Details. In our experiments, we set the number of encoder layers to 1 and the number of decoder layers to 4. Following previous work (Feng et al., 2022; Wang et al., 2024b), we adopt DIN as our sequential recommendation model and use pretrained Llama2-7b as our semantic encoder, with the hidden size set to 128 as reported in (Rajput et al., 2023; Wang et al., 2024b). The number of clusters in the hierarchical k-means is set to 256. To train our model, we use the Adam optimizer with a learning rate of 0.001 and apply a warmup strategy for stable training. The batch size is set to 256. Notably, UNGER demonstrates robustness to the hyperparameters of the CKA and IKD tasks, attributed to their fast convergence. Consequently, the loss coefficients $\alpha$ and $\beta$ are both set to 1.

### v3

@@ 3.3.4. Joint optimization
3.3.5. Unified Codes
After the training is finished, following previous work (Wang et al., 2024c), we apply the hierarchical k-means algorithm through an iterative residual quantization process. As depicted in Algorithm 1, this method allows us to efficiently encode high-dimensional item embeddings into discrete hierarchical codes, while preserving as much information as possible at each stage of quantization. The core mathematical relationship at each layer is defined as:
(6) $\mathbf{r}_{i}^{l}=\mathbf{r}_{i}^{l-1}-\mathbf{C}_{i}^{l}$
where $\mathbf{r}_{i}^{0}=\mathbf{v}_{i}$ denotes the original embedding of item $i$, and $\mathbf{C}_{i}^{l}$ represents the centroid assigned to item $i$ at the $l$-th layer of quantization. This recursive structure ensures that each subsequent residual $\mathbf{r}_{i}^{l}$ captures the portion of the embedding not explained by the current centroid approximation.
Algorithm 1 Hierarchical K-means Clustering
Input: Item embeddings $\mathcal{V}=\{\mathbf{v}_{1},\mathbf{v}_{2},\dots\}$, number of clusters $K$, hierarchy depth $L$
Output: Unicode sequences $\mathbf{c}_{i}=[c_{i}^{1},c_{i}^{2},\dots,c_{i}^{L}]$ for all items
1:for each item $i$ do
2:  Initialize residual: $\mathbf{r}_{i}^{0}\leftarrow\mathbf{v}_{i}$
3:end for
4:for layer $l\leftarrow 1$ to $L$ do
5:  Collect residuals: $\mathcal{R}^{l-1}\leftarrow\{\mathbf{r}_{1}^{l-1},\mathbf{r}_{2}^{l-1},\dots\}$
6:  Perform k-means clustering on $\mathcal{R}^{l-1}$ with $K$ clusters
7:  Store centroids in codebook $C_{l}=\{\mathbf{C}_{1}^{l},\mathbf{C}_{2}^{l},\dots,\mathbf{C}_{K}^{l}\}$
8:  for each item $i$ do
9:   Compute distances: $d_{k}=\|\mathbf{r}_{i}^{l-1}-\mathbf{C}_{k}^{l}\|^{2}$ for all $k\in[1,K]$
10:   Assign cluster index: $c_{i}^{l}\leftarrow\arg\min_{k}d_{k}$
11:   Record centroid: $\mathbf{C}_{i}^{l}\leftarrow\mathbf{C}_{c_{i}^{l}}^{l}$
12:   Update residual: $\mathbf{r}_{i}^{l}\leftarrow\mathbf{r}_{i}^{l-1}-\mathbf{C}_{i}^{l}$
13:  end for
14:end for
The algorithm proceeds as follows: We begin by initializing the residual vectors with the original item embeddings, i.e., $\mathbf{r}_{i}^{0}=\mathbf{v}_{i}$. At each hierarchical layer $l$, we collect all current residuals $\mathcal{R}^{l-1}$ and apply k-means clustering to partition them into $K$ clusters. This step identifies common patterns or structures in the embedding space that can be captured by a shared centroid.
The set of $K$ resulting centroids is stored as the codebook $C_{l}=\{\mathbf{C}_{1}^{l},\mathbf{C}_{2}^{l},\dots,\mathbf{C}_{K}^{l}\}$. For every item $i$, we compute the squared Euclidean distance between its current residual $\mathbf{r}_{i}^{l-1}$ and each centroid in $C_{l}$:
(7) $d_{k}=\|\mathbf{r}_{i}^{l-1}-\mathbf{C}_{k}^{l}\|^{2}$
The nearest centroid is then selected based on the minimum distance criterion:
(8) $c_{i}^{l}=\arg\min_{k}d_{k}$
Once the closest centroid $\mathbf{C}_{i}^{l}=\mathbf{C}_{c_{i}^{l}}^{l}$ is determined, it is subtracted from the current residual to compute the new residual for the next layer:
(9) $\mathbf{r}_{i}^{l}=\mathbf{r}_{i}^{l-1}-\mathbf{C}_{i}^{l}$
This subtraction removes the information captured by the current centroid, allowing the next layer to focus on the remaining details. This process is repeated iteratively across $L$ layers, with each layer progressively refining the approximation of the original embedding by capturing finer-grained residuals.
After all $L$ layers have been processed, we obtain a sequence of discrete cluster indices $\mathbf{c}_{i}=[c_{i}^{1},c_{i}^{2},\dots,c_{i}^{L}]$ for each item. This sequence serves as a compact and hierarchical unicode representation of the item embedding. The hierarchical structure allows for efficient indexing and retrieval while preserving high fidelity in the quantization process, as each level incrementally refines the representation by focusing on residuals not explained by previous levels.
Subsequently, we construct an item-unicode lookup table that maps each item to its corresponding unicode sequence. These unicodes serve as discrete identifiers for the items and play a crucial role in Stage II of our framework, where they are used to facilitate generative recommendation. By operating in the unicode space, the model can manipulate item representations in a structured and semantically meaningful manner, enabling flexible and efficient recommendation generation.
@@ 3.4. Stage II: Generative Recommendation
3.4.1. Encoding process
On the encoder side, given a user’s interaction history $X=[x_{1},x_{2},\dots,x_{t-1}]$, we input it into an encoder consisting of stacked multi-head self-attention layers and feed-forward layers, following the Transformer architecture. The encoder processes $X$ to produce a feature representation $H$, which captures the user’s interests and will be passed to the decoder.
3.4.2. Decoding process
On the decoder side, items are mapped to their unicodes derived in Stage I. Our objective changes from predicting the next item $x_{t}$ to predicting its item unicode $c_{1},c_{2},\dots,c_{L}]$. For training, a special token <BOS> is prepended to the item unicode to construct the decoder input. The generative recommendation loss is computed using a cross-entropy loss function, defined as follows:
@@ 3.4.2. Decoding process
3.4.3. Intra-modality knowledge distillation task
To compensate for the information loss caused by the quantization process in the first stage, we introduce an intra-modality knowledge distillation task to distill the global knowledge. Inspired by the use of the [CLS] token in BERT for capturing global context in classification tasks, we append a learnable token $c_{\text{dis}}]$ to the end of the decoder input, allowing gradients to propagate through each preceding unicode. This design encourages $c_{\text{dis}}]$ to capture global information.
Specifically, the final layer output corresponding to $c_{\text{dis}}]$ is used in a global contrastive learning objective over the item corpus: the positive sample is the integrated embedding $E_{t}$ of item $x_{t}$ learned from the first stage, and the negative sample is a randomly selected integrated embedding $E_{\text{neg}}$ from other items in the corpus, excluding $x_{t}$. This objective pulls the $c_{\text{dis}}]$ final layer output closer to the positive sample $E_{t}$ and pushes it away from the negative sample $E_{\text{neg}}$. The loss function for the distillation task is defined as follows:
@@ 3.4.3. Intra-modality knowledge distillation task
where $c_{\text{dis}}$ denotes the final layer output of the decoder for the special token $c_{\text{dis}}]$; $E_{\text{neg}}$ is the integrated embedding for a negative sample as learned in Stage I, while $E_{t}$ represents the integrated embedding of the target item, also learned in Stage I.
3.4.4. Joint optimization
Finally, in Stage II, we jointly optimize these two loss functions. The overall loss function for Stage II is defined as follows:
@@ 3.5. Training and Inference
3.5.1. Training
[…]
3.5.2. Inference
During inference phase, we rely exclusively on the Transformer model. In the decoding process, beam search is applied iteratively to generate each token within the item unicode sequence. Once finished, the item unicodes are mapped to their corresponding items through the item-unicode lookup table, creating a ranked recommendation list based on confidence scores to produce the final top-$k$ results.
@@ 3.6. Computational and Storage Costs Analysis
In this section, we analyze the computational and storage efficiency of our proposed method compared to existing approaches that handle multiple modalities separately. Our method leverages a unified code to jointly encode both semantic and collaborative information, which offers significant advantages in terms of both computational complexity and memory usage.
3.6.1. Computational Cost
In our framework, the encoding phase processes the user interaction sequence only once through the encoder, which yields a latent representation. This one-time encoding step ensures that the computational overhead during encoding remains minimal. The main computational burden lies in the decoding stage, where the decoder autoregressively generates each token of the target item code. Assuming the code length is fixed and the decoding of each item code takes $\mathcal{O}(n)$ time, the total decoding complexity for generating a recommendation is $\mathcal{O}(n)$.
In contrast, existing methods such as (Wang et al., 2024c; Kim et al., 2024) adopt separate modality-specific codes, typically maintaining $K$ parallel decoders—each corresponding to a different modality (e.g., semantic, collaborative, etc.). To generate recommendations, these methods must decode $K$ modality-specific codes in parallel, leading to a cumulative decoding cost of $\mathcal{O}(Kn)$. Moreover, these approaches usually include an additional ranking or fusion step based on confidence scores to integrate the outputs from different modalities. While we exclude the computational cost of this post-decoding ranking step in our comparison, it is worth noting that it may further increase the overall latency during inference.
In summary, by unifying all modalities into a single code and decoding only once, our approach significantly reduces the computational complexity from $\mathcal{O}(Kn)$ to $\mathcal{O}(n)$, offering a more scalable solution especially when the number of modalities $K$ is large.
3.6.2. Storage Cost
Let us now consider the memory requirements for storing the item codes. Suppose that storing one modality-specific code requires $\mathcal{O}(m)$ space. For methods that encode each modality independently, storing $K$ codes per item entails a total storage cost of $\mathcal{O}(Km)$. By contrast, our method compresses all relevant information into a single unified code, thus requiring only $\mathcal{O}(m)$ storage per item. This compression not only reduces the memory footprint but also simplifies the storage architecture and retrieval process during inference.
Overall, our unified coding strategy delivers substantial improvements in both computational and storage efficiency. The reduction from $\mathcal{O}(Kn)$ to $\mathcal{O}(n)$ in decoding time, and from $\mathcal{O}(Km)$ to $\mathcal{O}(m)$ in storage requirements, makes our approach well-suited for large-scale recommender systems where real-time performance and memory scalability are critical.
@@ 4.1. Experimental Setting
4.1.1. Dataset
We conduct experiments on three public benchmarks commonly used in the sequential recommendation task. For all datasets, we group the interaction records by users and sort them by the interaction timestamps ascendingly. Following (Rendle et al., 2010; Zhang et al., 2019), we only keep the 5-core dataset, which filters unpopular items and inactive users with fewer than five interaction records. Statistics of these datasets are shown in Table 2.
@@ 4.1.1. Dataset
Table 3. Performance comparison of different methods. The best performance is highlighted in bold while the second best performance is underlined. All the results of UNGER are statistically significant with paired t-test $p<0.05$ compared to the best baseline models. We computed the standard error for all metrics for the datasets by running the training and evaluation 5 times using different random seeds chosen in [2020, 2021, 2022, 2023, 2024].
Dataset Metric Classical Generative UNGER (Ours)
Beauty Recall@10 0.0283 0.0347 0.0605 0.0347 0.0512 0.0407 0.0647 0.0664 0.0617* 0.0524* 0.0836 0.0939 $\pm$ 0.00093
Recall@20 0.0479 0.0556 0.0902 0.0599 0.0773 0.0656 0.0994 0.0915 0.0924* 0.0807* 0.1124 0.1289 $\pm$ 0.00084
NDCG@10 0.0137 0.0176 0.0318 0.0170 0.0266 0.0208 0.0327 0.0400 0.0339* 0.0263* 0.0525 0.0559 $\pm$ 0.00037
NDCG@20 0.0187 0.0229 0.0394 0.0233 0.0332 0.0270 0.0414 0.0464 0.0417* 0.0335* 0.0599 0.0646 $\pm$ 0.00041
Sports Recall@10 0.0204 0.0194 0.0350 0.0191 0.0313 0.0288 0.0385 0.0247 0.0376* 0.0348* 0.0441 0.0471 $\pm$ 0.00078
Recall@20 0.0333 0.0314 0.0507 0.0315 0.0477 0.0463 0.0607 0.0375 0.0577* 0.0533* 0.0659 0.0710 $\pm$ 0.00075
NDCG@10 0.0110 0.0097 0.0192 0.0099 0.0159 0.0156 0.0204 0.0133 0.0196* 0.0179* 0.0236 0.0259 $\pm$ 0.00034
NDCG@20 0.0142 0.0126 0.0231 0.0130 0.0201 0.0200 0.0260 0.0164 0.0246* 0.0226* 0.0291 0.0319 $\pm$ 0.00049
Toys Recall@10 0.0176 0.0270 0.0675 0.0203 0.0497 0.0381 0.0700 0.0383 0.0578* 0.0474* 0.0714 0.0822 $\pm$ 0.00085
Recall@20 0.0301 0.0420 0.0941 0.0358 0.0716 0.0632 0.1065 0.0483 0.0838* 0.0704* 0.1024 0.1154 $\pm$ 0.00070
NDCG@10 0.0084 0.0141 0.0374 0.0099 0.0277 0.0189 0.0376 0.0285 0.0321* 0.0242* 0.0489 0.0505 $\pm$ 0.00032
NDCG@20 0.0116 0.0179 0.0441 0.0138 0.0332 0.0252 0.0468 0.0310 0.0386* 0.0300* 0.0538 0.0573 $\pm$ 0.00025
4.1.2. Evaluation Metrics
@@
We employ two broadly used criteria for the matching phase, i.e., Recall and Normalized Discounted Cumulative Gain (NDCG). We report metrics computed on the top 10/20 recommended candidates. The used evaluation metrics are as follows:
  • Recall@$K$: It measures the proportion of correctly predicted relevant items in the top-$K$ recommendation list:
    (13) $\text{Recall@}K=\frac{1}{|\mathcal{U}|}\sum_{u\in\mathcal{U}}\mathbb{I}(r_{u}<K),$
    where $\mathcal{U}$ is the set of users in the test set, and $r_{u}$ denotes the rank position of the ground-truth item $i_{u}$ in the top-$K$ recommendation list for user $u$.
  • NDCG@$K$: It evaluates the ranking quality of the top-$K$ recommended items by considering both item relevance and order:
    (14) $\text{DCG}_{u}@K=\sum_{j=1}^{K}\frac{2^{y_{u,j}}-1}{\log_{2}(j+1)},$
    (15) $\text{iDCG}_{u}@K=\sum_{j=1}^{K}\frac{2^{y^{*}_{u,j}}-1}{\log_{2}(j+1)},$
    (16) $\text{NDCG}_{u}@K=\frac{\text{DCG}_{u}@K}{\text{iDCG}_{u}@K},$
    (17) $\text{NDCG@}K=\frac{1}{|\mathcal{U}|}\sum_{u=1}^{|\mathcal{U}|}\text{NDCG}_{u}@K,$
    where $y_{u,j}$ denotes the relevance of the $j$-th recommended item for user $u$, and $y^{*}_{u,j}$ is the relevance in the ideal ranking. In binary relevance settings, $y_{u,j}=1$ if the $j$-th item is the ground-truth item, and 0 otherwise. The final NDCG@$K$ is the average over all users.
Following the standard evaluation protocol (Kang and McAuley, 2018), we use the leave-one-out strategy for evaluation. For each item sequence, the last item is used for testing, the item before the last is used for validation, and the rest is used for training. During training, we limit the number of items in a user’s history to 20.
4.1.3. Implementation Details
@@
In our experiments, we set the number of encoder layers to 1 and the number of decoder layers to 4. Following previous work (Feng et al., 2022; Wang et al., 2024c), we adopt DIN as our sequential recommendation model and use pretrained Llama2-7b as our semantic encoder, with the hidden size set to 128 as reported in (Rajput et al., 2023; Wang et al., 2024c). The number of clusters in the hierarchical k-means is set to 256. To train our model, we use the Adam optimizer with a learning rate of 0.001 and apply a warmup strategy for stable training. The batch size is set to 256. Notably, UNGER demonstrates robustness to the hyperparameters of the CKA and IKD tasks, attributed to their fast convergence. Consequently, the loss coefficients $\alpha$ and $\beta$ are both set to 1. Each experiment is conducted five times and the average score is reported. All model parameters and settings are shown in Table 4.
Table 4. Model hyperparameters and settings.
Parameter Value
Embedding Dimension 96
Model Layers 4
Hidden Size 256
Heads 6
Num of Clusters 256
Cluster Depth 4
Learning Rate 1e-3
Optimizer Adam
Semantic Encoder Llama2-7b
Collaborative Encoder DIN
Batch Size 256
Training Steps 20000
Dropout Rate 0
Beam Width 100
Activation Function ReLU
Weight Decay 1e-7
Warmup Steps 2000
Warmup Initial Learning Rate 1e-7
Random Seed 2024
$\alpha$ 1.0
$\beta$ 1.0
$\tau$ 1.0
Table 5. The information about practical resources required.
Information Value
Device $\text{NVIDIA RTX 3090 GPU (24GB)}\times 1$
Training Time less than 1 GPU hour
Inference Latency (ms) 19.5
Model Parameter 10249988